### v1

# Multimodal MRI-based Detection of Amyloid Status in Alzheimer’s Disease Continuum

Abstract
Amyloid- (A) plaques in conjunction with hyperphosphorylated tau proteins in the form of neurofibrillary tangles are the two neuropathological hallmarks of Alzheimer’s disease (AD). In particular, the accumulation of A plaques, as evinced by the A/T/N (amyloid/tau/neurodegeneration) framework, marks the initial stage. Thus, the identification of individuals with A positivity could enable early diagnosis and potentially lead to more effective interventions. Deep learning methods relying mainly on amyloid PET images have been employed to this end. However, PET imaging has some disadvantages, including the need of radiotracers and expensive acquisitions. Hence, in this work, we propose a novel multimodal approach that integrates information from structural, functional, and diffusion MRI data to discriminate A status in the AD continuum. Our method achieved an accuracy of . Furthermore, a post-hoc explainability analysis (guided backpropagation) was performed to retrieve the brain regions that most influenced the model predictions. This analysis identified some key regions that were common across modalities, some of which were well-established AD-discriminative biomarkers and related to A deposition, such as the hippocampus, thalamus, precuneus, and cingulate gyrus. Hence, our study demonstrates the potential viability of MRI-based characterization of A status, paving the way for further research in this domain.
1 Introduction
The amyloid cascade hypothesis in Alzheimer’s Disease (AD) posits that the accumulation of extracellular Amyloid- (A) neuritic plaques in the brain leads to tauopathy and consequent neurodegeneration ([1]). Hence, A deposition in the brain is considered to be the first step and the principal trigger of AD pathology ([2, 1]). In consequence, neurodegeneration of gray matter (GM) may be related to the deposition of A plaques, resulting in cerebral atrophy and synaptic loss ([3]), altering many brain regions, especially subcortical areas, and functional networks. Additionally, A plaques in AD are also linked with demyelination of white matter (WM) tracts ([4]).
The precise identification of A proteins, and thus the classification of patients as A positive or negative, is only possible through amyloid positron emission tomography (PET) scans and cerebrospinal fluid (CSF) tests. Despite their undeniable utility, their use has some negative implications. Specifically, amyloid PET uses ionizing radiation and radioactive tracers, is expensive, and is not widely available ([5, 6]). Additionally, the CSF test requires an invasive lumbar puncture.
In contrast, common, non-invasive MRI techniques can help identifying the effects of A plaques in the brain. In particular, structural MRI (sMRI) has been widely used for AD detection and early-prediction due to its ability to detect GM atrophy and structural changes. Resting-state functional MRI (rs-fMRI) detects changes in blood oxygenation level-dependent (BOLD) signals, which depend on neurovascular coupling, and hence indirectly measures brain neural activity ([7]). Different resting-state networks (RSNs), like the saliency and default-mode (DM) networks, are closely related to AD pathology, showing functional alterations ([8, 9, 10]). Finally, diffusion MRI (dMRI) is an imaging technique that relies on the movement of water molecules to mark WM boundaries enabling the reconstruction of fiber trajectories and the study of brain structural connections in AD ([4]). Additionally, dMRI data are used to extract microstructural information ([11]). These two modalities are generally used to map whole-brain connectomes, with rs-fMRI describing the functional connectivity (FC) between regions (usually in terms of correlations) and dMRI the structural connectivity (SC) (most often relying on the number of WM fibers).
Currently, different studies focusing on A classification tasks (positive vs negative) in AD research have employed PET images and related features with deep learning (DL) models, while only a few works used MRI data. [12] considered a slice-level approach for the identification of the A status in conjunction with a 2D-CNN for feature extraction and classification. In the same manner, [13] adopted the slice-level approach for the classification of A status employing two different 2D-CNN (ResNet-50 and ResNet-152). These two works were able to reach a high level of accuracy (around ) using a slice-level method. On the other hand, [6] considered three different well-known 3D-CNNs for detecting A positivity relying on 3D Florbetaben brain PET images. [14] developed a particular CNN composed of different submodules for analyzing the 3D FDG-PET images, converting them into slices following the three different axes. [15] faced the problem of A classification relying on both early and delay-phase FBB PET images and tested their models considering both single and fused modalities. Using FDG and Amyloid PET-derived images, they were able to reach competitive performance in this task, with accuracies around . More recently, [16] employed FDG-PET-derived features in order to predict the A status in the AD continuum. To this end, they tested eight different feature selection methods and eight different classifiers. LASSO in conjunction with the Gaussian Naive Bayes (GNB) model performed better with respect to the others, achieving an AUC of 0.924. Regarding MRI-based approaches for A detection, [17] employed sMRI images in conjunction with a 3D-CNN, while [18] used an SDF-based convolutional network to analyze the hippocampus region. Using sMRI-derived images and features, they were able to reach accuracies around . Due to the heterogeneous factors that lead to AD, in recent years, many studies have focused on multimodal DL models due to their ability to integrate information of different nature and to outperform single-modality methods ([19, 17, 20]). Moreover, among the different DL models, graph neural and convolutional networks (GNNs/GCNs) are becoming popular in neuroscience due to their perfect fit to functional and structural brain networks. [21] proposed a benchmark for analyzing fMRI and dMRI networks through GNNs, testing different messages passing, node features, and pooling operations, while Wee and colleagues developed a GCN to study the cortical thickness ([22]).
Although DL models can achieve high performance in different tasks, they do not easily provide interpretable output for what they have learned, which is particularly problematic in clinical and biomedical domains. To address this issue, eXplainable Artificial Intelligence (XAI) methods have been developed allowing to identify the contributions of input features to final predictions potentially highlighting crucial physiological information for AD ([23, 24, 25, 26]).
In this study, we present a multimodal and explainable DL framework for the classification of A positive and negative status in the AD continuum. Our framework included structural and functional connectomes derived from dMRI and rs-fMRI, respectively, along with sMRI-derived GM 3D volumes in order to investigate complementary aspects as well as their relations. This was complemented with an extensive post-hoc XAI analysis to uncover the input features that most influenced the final predictions.
2 Methods
2.1 Dataset
The sMRI, rs-fMRI, and dMRI neuroimaging data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial MRI, PET, other biological markers, and clinical and neuropsychological assessment could be combined to measure the progression of mild cognitive impairment (MCI) and early AD. For up-to-date information, please refer to www.adni-info.org.
One of the key strengths of the dataset is the inclusion/exclusion criteria adopted to recruit the subjects; subjects with neurological diseases other than AD and with different substance/drug use were excluded from the study. Due to this, the results we uncover related to A are unlikely to be related to other diseases or the use of a particular substance. For more information about exclusion criteria, please refer to the official document at this link: https://adni.loni.usc.edu/wp-content/themes/freshnews-dev-v2/documents/consentForms/ADNI3_ProtocolVersion3.1_20201204.pdf.
In this work, our dataset was initially composed of preprocessed sMRI images (from subjects) from the ADNI , , , and GO phases, out of which passed quality control (QC) (from subjects). For rs-fMRI, preprocessed images were considered (from individuals) from the ADNI , , and GO phases, out of which passed QC (from individuals). For dMRI, preprocessed images (from subjects) from ADNI were considered, out of which (from subjects) passed QC. Additional information about QC is detailed in the next paragraph. The images of the first available timepoint from only subjects belonging to control (CN), significant memory concern (SMC), early MCI (EMCI), late MCI (LMCI), and AD clinical classes, that had all three modalities, and available A status were included. Lumbar puncture to retrieve CSF samples was performed using the procedures described on the ADNI website, and subjects were labeled as A positive or negative based on the A protein levels reported by the CSF test. Similarly to [5], a cutoff of was used to define the A status (i.e., for positivity). After that, A negative CN, SMC, and EMCI (, , and subjects, respectively) were considered as ’control class’, while A positive EMCI, LMCI, and AD (, , and patients, respectively) formed the ’patient class’, resulting in a final cohort of individuals. Table 1 shows the demographic information of the final cohort.
Status # of subjects Age Sex (M/F) MMSE A CN SMC EMCI LMCI AD A - - - A + - -
The MRI images for the considered cohort were collected as follows: i) T1-weighted sMRI: TR/TE=shortest, TI= ms, FOV= mm2, mm isotropic resolution, slices=; ii) rs-fMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, volumes in almost all subjects, with minimal variations (e.g., ) in a small subset; iii) single-shell dMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, b= and s/mm2.
2.2 Preprocessing and Feature extraction
The sMRI preprocessing included tissue segmentation of GM, WM, and CSF with the modulated normalization algorithm in the statistical parametric mapping toolbox (SPM12, http://www.fil.ion.ucl.ac.uk/spm/). This work used GM volumes smoothed with a Gaussian kernel (FWHM=mm). For QC, images that had a low correlation with individual and/or group-level masks were discarded, which involved correlating data at three levels: the entire image, the top slices, and the bottom slices. The full preprocessed GM volume was input to the sMRI channel of the neural network, resulting in an input size of for each subject.
Adhering to the process proposed in [27], the rs-fMRI data was preprocessed with SPM12 including rigid body motion correction, removal of scans with high head motion parameters ( of rotations and in translations), slice-timing correction, warping to the standard MNI space using the EPI template, resampling to isotropic voxels, and smoothing with a Gaussian kernel (FWHM = mm). QC was the same as for sMRI, correlating the data at three levels: the entire image, the top slices, and the bottom slices. Fifty-three maximally independent components (ICs) covering the whole brain were extracted using spatially constrained ICA with the Neuromark_fMRI_ template (available in the GIFT software; http://trendscenter.org/software/gift). The ICs were divided into RSNs: the i) Sub-cortical (SuC); ii) Auditory (AU); iii) Sensorimotor (SM); iv) Visual (VI); v) Cognitive-control (CC); vi) DM); and, vii) Cerebellar (CB) networks. For each subject, the Pearson correlation between IC time courses was computed, resulting in a x static functional network connectivity (FNC) matrix, where FNC is the network analog of FC in that the timecourses represent weighted partially overlapping whole brain patterns. Finally, each FNC matrix was converted into a complete, undirected, and weighted graph. The edges’ weights correspond to the FNC correlation values, considering both positive and negative values, while the values of the nodes (i.e., ICs) were initialized at a value of one in order to force the network to learn a latent representation based only on the connectivity information. This FNC-based graph was the input to the rs-fMRI channel.
The dMRI volumes were preprocessed via brain extraction followed by Eddy currents correction (FSL 6.0, https://fsl.fmrib.ox.ac.uk/). The data was then denoised using local principal component analysis (PCA) via empirical thresholds relying on the Python dipy library. Subsequently, nonlinear registration to the MNI space was applied to correct for EPI-induced currents. QC was performed during preprocessing by visual inspection of images before and after registration. MRtrix (https://www.mrtrix.org/) was used to derive an anatomically constrained probabilistic tractography ( million streamlines, step=mm, maximum length=mm, and backtracking) filtered with SIFT2 ([28]). Subject-specific brain parcellations from T1-weighted images were derived using FreeSurfer (https://surfer.nmr.mgh.harvard.edu/) and used as regions of interest (ROIs) in the SC calculation. The SC matrix was built by counting the number of streamlines connecting all pairs of regions from the Desikan-Killiany ([29]) structural atlas, ignoring self-connections. Similarly to rs-fMRI, each SC matrix was converted into a complete, undirected, and weighted graph. In this case, the edges’ weights were defined as the number of streamlines between pairs of ROIs, and the values of the nodes (i.e., anatomical ROIs) were initialized to one. This SC-based graph was the input to the dMRI channel.
2.3 Framework architecture
The proposed framework is shown in Figure 1. In detail, the DL architecture used for the classification of A status (positive/negative) has two modules: i) a feature reduction module that actuates feature reduction using three different Neural Networks (NNs) to transform the input data into corresponding latent representations; and ii) a data fusion & classification module that concatenates the latent representations of each modality and uses a Multilayer Perceptron (MLP) that takes the fused latent features as input for the final classification. Finally, a post-hoc explainability analysis was performed on the correctly classified subjects to obtain feature contributions for the classification task. In the following paragraphs, the two modules are further described.
Feature reduction module
Three different NNs form the feature reduction module. Each NN extracts a latent representation of one modality resulting in a latent vector of features. The 3D sMRI volumes were analyzed using a 3D-CNN with six convolutional layers and three max-pooling layers, completed with four dense layers. Conversely, both FNC and SC graphs were analyzed using two different GCNs. These models update the representation of the node aggregating neighbor information, iteratively in each layer, through the message-passing scheme. The rs-fMRI channel was analyzed using a GCN with two graph convolutional layers proposed by [30] that were followed by a dense layer. This type of convolutional layer is a powerful operator that integrates both high- and low-level structures along with their relationships into a single graph ([30]). The dMRI channel was analyzed with a different GCN with two Chebyshev spectral graph convolutional layers proposed by [31] and then followed by a dense layer. Spectral convolutional layers are high-performance layers that have been used effectively in irregular graphs ([32]). These two architectures were chosen on an empirical basis as those leading to the best unimodal classification accuracy.
Data fusion & Classification module
The data fusion & classification module consists of a fusion layer and a classifier. The fusion layer concatenates the latent features extracted from the three channels, resulting in a vector of features that incorporate information from all three modalities for each subject. Lastly, the latent vector was used as input for the final MLP classifier, composed of three dense layers. The convolutional and dense layers in this framework used a ReLU activation function, except for the last layer, which used a softmax activation function to obtain the classification probabilities for each class.
2.4 Training scheme and evaluation
The model was trained with stratified -fold cross-validation on the entire cohort to investigate its generalizability across individuals, and a hyperparameter search was performed empirically. The mini-batch strategy (with subjects per batch) was adopted. The Adam optimizer (learning rate: ) was used to update the entire multimodal architecture. L regularization (weight decay: ) was applied to reduce overfitting. Weighted cross entropy was used as the loss function. The model was trained for epochs.
Performance was evaluated using the mean evaluation accuracy, precision, recall, and F1 score over the five folds.
For comparison, we also tested the unimodal models with the corresponding MRI data, where the architectures were the same of the different branches of the multimodal framework (i.e., 3D-CNN for sMRI, and GCN for both rs-fMRI and dMRI).
3 Post-hoc analysis
3.1 Guided backpropagation
The post-hoc XAI analysis was conducted using guided backpropagation (GBP) ([33]). GBP uses the model gradients to extract the feature contribution maps with the same shape as the input data. It belongs to the "modified backpropagation" class of XAI methods in which the backward flow of gradients is modified with ReLU activation ([34]), setting the negative gradients to zero and only allowing non-negative gradients to be backpropagated. This approach enables the visualization of which input features activated the neurons and most contributed to the final prediction.
3.2 Contribution maps and Statistical analysis
The attribution maps were extracted for the correctly classified A positive subjects, and in particular, we focused on the average A positive attribution map to identify the most important features. To evaluate the sMRI GBP contribution maps, the Harvard-Oxford ([29]) and the probabilistic cerebellar ([35]) atlases from FSL were employed to define different ROIs , including cortical, subcortical, and cerebellum regions. The sum of GBP attributions inside each ROI was calculated for the sMRI and weighted to account for the volume of each specific region. Conversely, the GBP attribution of each node for both rs-fMRI and dMRI was extracted directly from the two GCNs. Then, considering the average map for each modality, the percentage of explanation for each region/node was computed over the total contribution within and between each modality. The top ROIs (sMRI) and nodes (rs-fMRI and dMRI) with the highest percentage of GBP contribution were selected for further investigation.
Subsequent statistical analyses were performed on the original data for all correctly classified subjects. For the sMRI, the mean values of the top ROIs resulting from the XAI analysis were extracted from the input GM volumes and used as features for the statistical analysis. For the rs-fMRI and dMRI, graph-based measures were first derived from the full connectivity matrices in order to have a summary measure per node, and then we considered only the top nodes for both rs-fMRI and dMRI for the statistical analysis. In particular, the node strength was computed for rs-fMRI and FNC matrices. This is defined as the sum of the weights of the edges connected with a given node, where in absolute terms, higher values mean more important nodes. Betweenness centrality was calculated for SC matrices, representing the fraction of all shortest paths in the SC matrix that contains the node under analysis, where the shortest path is the shortest sequence of nodes between node i and node j. As the sparsity of the SC matrix could limit the interpretation of the node strength results for the dMRI data, we preferred to rely on a centrality measure for this analysis. In this case, nodes that belong to more paths likely play a pivotal role in the propagation of the information inside the network. These two metrics were computed using the Brain Connectivity Toolbox (BCT) ([36]) in Matlab.
Mann-Whitney tests were then used to compare the values of all these features for the top ROIs/nodes between A positive and negative correctly classified patients. Finally, FDR correction for multiple comparisons was applied to the p-values obtained from the statistical tests.
4 Results
4.1 Classification performance
The proposed framework for the classification of A status achieved a meanstd accuracy, precision, recall, and F1 score of , , , and , respectively, across the evaluation folds.
The single networks that composed the multimodal framework along with the corresponding input data were also tested in the same classification task. Table 2 shows the performance comparisons for the multimodal method and the three unimodal models. Results highlight how the multimodal framework is able to outperform the unimodal models in terms of accuracy, recall, and F1 score in the same classification task.
Model ACC PRE REC F1 Multimodal Unimodal sMRI Unimodal fMRI Unimodal dMRI
4.2 GBP-based attribution maps
The evaluation set in the fold with the highest evaluation accuracy was used for the post-hoc analysis. Importantly, in the analyzed fold, the evaluation subjects were not used for training.
Figure 2A shows the sMRI GBP attribution map for the mean A positive patient overlaid to the MNI template. Qualitatively, mainly subcortical regions (e.g., hippocampus, thalamus) are relevant to the final classification along with a few cortical areas.
Figure 2B shows the most important nodes for the rs-fMRI modality. All displayed nodes (ICs) are bihemispheric, except for the left inferior parietal lobule (IC38) node, which is mainly in the left hemisphere. The most important nodes belong to the DM ( nodes), CC ( nodes), VI ( nodes), and SM ( node) networks.
Finally, Figure 2C shows the most important dMRI nodes (ROIs). Subcortical regions (thalamus, cerebellum) along with areas in the frontal, temporal, and parietal lobes are identified.
4.3 Feature relevance & Statistical analysis
Figure 3A shows violinplots for the top sMRI features, while Table 3a shows the corresponding percentages of mean GBP contribution across A positive correctly classified patients for the same regions, weighted by their volumes. The Table also reports the p-values and FDR-corrected p-values of the statistical tests performed on the input features. We identified significant differences between classes after FDR correction in the precuneus, cingulate gyrus (posterior division), thalamus, hippocampus, supracalcarine and intracalcarine cortices, amygdala, and temporal occipital fusiform cortex showing an overall comparison direction of .
Figure 3B shows violinplots representing the graph node strength for the top rs-fMRI nodes, while Table 3b shows the mean percentages of GBP contribution extracted from the same nodes (ICs) across A positive patients. After FDR correction, the precuneus (IC48) was the only region with significant differences (). Before correction, another precuneus component (IC43) and precentral gyrus (IC14) were also significant with , while the left inferior parietal lobule (IC38) exhibited a trend towards uncorrected statistical significance (p, with ).
Figure 3C shows violinplots representing the graph betweenness centrality for the top dMRI nodes, and Table 3c shows the mean percentages of GBP contribution for the same nodes across A positive correctly classified patients. The precuneus LH and cerebellum RH were significant when considering uncorrected p-values (), but none of the results survived multiple comparison corrections. Moreover, the right superior parietal gyrus showed a trend towards significance (p, ).
5 Discussion
In this work, we presented a multimodal data fusion framework that integrates multiple MRI techniques (sMRI, rs-fMRI, and dMRI) for the classification of A status. A post-hoc XAI analysis and evaluation was performed using GBP to identify feature importance, complemented by statistical analyses on the identified input features for validation via post-hoc plausibility assessment. The underlying assumption is that maximally contributing regions would bring discriminative information across the target classes.
a) Structural MRI
Region
Percentage (%)
p-value
FDR adj.
p-value
Comparison
direction
Accumbens† (Acc)
n.s.
Cingulate gyrus, posterior† (CGp)
0.00182
0.00304
NEGPOS
Putamen† (Put)
n.s.
Thalamus† (Thl)
0.00016
0.00146
NEGPOS
Hippocampus† (Hipp)
0.00073
0.00146
NEGPOS
Supracalcarine cortex† (SccC)
0.00043
0.00146
NEGPOS
Amygdala† (Amy)
0.00066
0.00146
NEGPOS
Temporal occipital fusiform cortex† (TOF)
0.01793
0.02561
NEGPOS
Precuneus cortex† (PcC)
0.00066
0.00146
NEGPOS
Intracalcarine cortex† (IcC)
0.03509
0.04386
NEGPOS
b) Resting-state functional MRI
Region
RSN
Percentage (%)
p-value
FDR adj.
p-value
Comparison
direction
Middle frontal gyrus∗ (IC31)
CC
NEGPOS
Calcarine gyrus (IC17)
VI
n.s.
Precentral gyrus (IC14)
SM
0.04965
NEGPOS
Precuneus∗ (IC48)
DM
0.00389
0.03892
NEGPOS
Left inferior parietal lobule (IC38)
CC
NEGPOS
Middle frontal gyrus∗ (IC36)
CC
n.s.
Hippocampus (IC37)
CC
n.s.
Precuneus∗ (IC43)
DM
0.02623
NEGPOS
Lingual gyrus (IC24)
VI
n.s.
Posterior cingulate cortex (IC46)
DM
n.s.
c) Diffusion MRI
Region Percentage (%) p-value FDR adj. p-value Comparison direction Rostral middle frontal LH (L.RMFG) n.s. Middle temporal LH (L.MTG) n.s. Inferior parietal RH (R.IPG) n.s. Precuneus LH (L.PCU) 0.01470 NEGPOS Cerebellum RH (R.CER) 0.03887 NEGPOS Cerebellum LH (L.CER) n.s. Lateral orbito frontal RH (R.LOFG) n.s. Superior parietal RH (R.SPG) NEGPOS Thalamus RH (R.TH) n.s. Supra marginal LH (L.SMG) n.s.
The term n.s. means not statistically significant, the sign † means top regions across the three modalities, while the sign ∗ means ICs with different spatial locations in rs-fMRI, and LH/RH means left/right hemisphere, respectively, in dMRI.
5.1 Classification performance
Our results showed that the multimodal framework outperformed single-modality models in terms of classification performance, in particular for rs-fMRI and dMRI, while the sMRI achieved performance close to the multimodal one, but with high variance across the evaluation folds. This confirms the added value provided by multimodality approaches which improve classification accuracy and, at the same time, carry more information, particularly relevant for the study of complex neurodegenerative diseases influenced by multi-domain factors.
Table 4 shows the performance of state-of-the-art works for the classification of A positive and negative subjects. Of note, current state-of-the-art works use multiple datasets and approaches for addressing this task without considering a common benchmark dataset. Hence, the comparison of the different approaches is imperfect and any resulting conclusions held tentatively. Most of the works relied on PET scans to address this task, e.g., [14] developed a CNN model that takes in input the different views (axial, coronal, and sagittal) of a 3D volume; while [6] used three different well-known 3D-CNN (Inception3D, ResNet3D, and VGG3D) to address this task using the full 3D volume of PET scans as input. They achieved an accuracy of around on average in the test set, with a maximum accuracy of by [6] using the VGG3D architecture. In the same way, [37] employed a 3D-CNN to identify the A status reporting both the average validation accuracy and also the accuracy on a hold-out test set (ADNI data) using the ensemble method created with the best models across the different folds, achieving accuracies of around . Recently, [16] employed different combinations of feature selection methods and classifiers in order to analyze FDG-PET-derived features. They achieved an AUC of employing LASSO with the GNB model. On the other hand, only few works have used MRI modalities to classify A positive and negative subjects. [18] used only the hippocampus region for predicting A positivity, testing their network on two different classification tasks: AD A POS versus CN A NEG, and MCI A POS versus MCI A NEG. They achieved an accuracy of and , in the first and second tasks, respectively. Lastly, [17] implemented a 3D-CNN for detecting A status from sMRI images defining the A status of the CN, MCI, and AD patients considering the PET cortical standardized uptake value ratio. They achieved an accuracy of .
Comparing our proposed method with these five works, we outperformed the method that relied on 3D FDG-PET proposed by [14] and the performance of [17]. Our study achieved competitive performance in the validation set compared to [6] using amyloid PET images, although our framework achieved a higher accuracy compared to the ResNet3D they used, and a similar accuracy to the one reached by the model proposed by [18] using the hippocampus region only. While the current classification of A positive or negative subjects can only be performed through CSF tests or amyloid PET scans with high precision, the development of different approaches based on non-invasive MRI modalities for determining A status could improve patient care highlighting specific mechanisms that underlie AD. We further showed the ability of our model to successfully integrate structural GM features extracted from a 3D-CNN along with functional and structural connectivity analyzed by two types of GCNs. Notably, while a few studies have recently started to explore SC/FC in conjunction with GCNs in the AD classification tasks ([38, 39]), results are still limited in this specific problem calling for further investigation in the AD continuum.
Authors Modalities Study cohort Input data Model ACC REC PRE F1 [14] A/FDG-PET POS, NEG 3D FDG-PET 2.5-D CNN - - n.d. - [6] A PET POS, NEG 3D A PET Inception3D ResNet3D VGG3D - - - - - - n.d. n.d. [37] A PET POS, NEG n.d. 3D A PET 3D-CNN - - n.d. - [16] FDG-PET POS, NEG FDG-PET-derived features LASSO+GNB (AUC) n.d. n.d. n.d. [18] sMRI AD POS, CN NEG MCI POS, MCI NEG Hippocampus region SDF-based NN n.d. n.d. n.d. [17] sMRI POS, NEG n.d. CN, MCI, AD 3D volume 3D-CNN n.d. n.d. Proposed framework sMRI, fMRI, dMRI POS, NEG GM volume FNC graph, SC graph Multimodal DL model
5.2 Explainability analysis
The post-hoc analysis, performed on the correctly classified subjects, consisted of two steps: i) extraction of GBP attribution maps for A positive patients and corresponding percentage of contribution for each brain region both within and between modalities for the average positive patient; ii) statistical analysis on the input features of the most important regions/nodes identified by GBP.
In our analysis of sMRI attribution maps, multiple brain regions involved in AD neurodegeneration were identified. These regions had a higher percentage of contribution relative to the others. These regions belonged to subcortical areas (hippocampus, thalamus, putamen, accumbens, and amygdala) and temporal/occipital/parietal areas (posterior cingulate gyrus, precuneus cortex, supracalcarine and intracalcarine cortices, and the temporal occipital fusiform cortex). The hippocampus, in particular, is a well-known biomarker for AD that is subject to high levels of atrophy. Some studies have suggested that this atrophy is attributable to the deposition of A plaques ([40]). An important region that is connected with the hippocampus is the cingulate gyrus which showed a strong reduction of GM in AD patients ([41, 42]). In a clinical study performed by [43], significant associations were detected between A accumulation and GM atrophy in the hippocampus and posterior cingulate gyrus for the MCI and AD A+ patients. Moreover, other subcortical structures, like the thalamus and putamen, have a high atrophy rate in clinical patients relative to healthy individuals ([44, 43]), probably due to A deposition. A previous study highlighted an increase in SUVR (derived from florbetapir PET amyloid) in both putamen and thalamus in the preclinical stages of AD ([45]). In the statistical analysis performed on the input GM volumes, significant differences were detected in the hippocampus, posterior cingulate gyrus, amygdala, thalamus, and precuneus, which demonstrated a considerable decrease of GM in the A positive patients. These findings were consistent with preexisting literature. Additionally, [46] highlighted other regions (that we also found to be relevant) involved in the different phases of A deposition (i.e., proisocortex, allocortical areas, diencephalic nuclei, and striatum). For the rs-fMRI, the brain regions (ICs) that most contributed to the final classification resided in the DM and CC networks, along with three regions for VI and SM networks (two and one, respectively). The DM network is involved in memory, self-knowledge, and thinking, and it is highly related to AD ([8, 41]) Previous studies also suggested that DMN is associated with A plaques ([47]). The CC network is associated with selective attention, working memory, and stimulus-response mapping ([48, 49]). The statistical analysis, performed on the node strength values, detected one region with significant differences between the two groups after FDR correction and three regions before (uncorrected p-values). One more region, the left inferior parietal lobule, exhibited a trend toward significance. Specifically, the A positive class had a significantly increased node strength in the precentral gyrus (IC14, SM network), which was consistent with the findings of [50] and [51]. The precuneus regions (IC43 and IC48, DM network) were also significantly different between groups with increased strength in the positive group. As in other works ([52, 53]), this suggested that the precuneus plays a key role and undergoes pathological changes related to AD. Previous works also identified the posterior DM (the precuneus and posterior cingulate cortex) ([9]) and parts of the CC network (i.e., the middle frontal gyrus and hippocampus) ([54, 55]) as strongly affected by A deposition. Lastly, in the dMRI channel, GBP identified multiple cortical and subcortical regions as relevant to the final predictions of our model. The cerebellum was also revealed to be important by the model. Of note, the rostral middle frontal gyrus was assigned a high percentage of contribution relative to the other regions. Additionally, the middle temporal gyrus, precuneus, thalamus, and superior parietal regions showed high importance relative to the other regions. After FDR correction, no region had statistically significant group differences in betweenness centrality values. On the other hand, based upon uncorrected p-values, significant differences (p) were highlighted in the precuneus and cerebellum (right hemisphere), while the superior parietal region was close to significance (p=). Regarding the two nodes with significant uncorrected differences, the precuneus, which is also identified in the rs-fMRI, is an important hub for functional operations because it is highly connected with other regions by both short- and long-range WM fibers ([53]). On the other hand, different studies highlighted how the cerebellum is subject to increase A deposition in AD pathology ([56, 57]) at different stages (stage in [58] and stage in [46]) relative to normally aged subjects. Widespread structural alterations in WM tracts connecting cortical and subcortical regions in both hemispheres were detected between A negative CN and positive preclinical AD patients by [59]. Finally, when combining the explanations of all channels, we found that sMRI most contributed to the final prediction. This was expected as we initially assumed that the effects of A plaques would predominantly impact GM volumes.
5.3 Common regions across modalities
Multiple brain regions were identified as important by GBP across all three channels. Specifically, the precuneus was highlighted in all three modalities. It was one of the most interesting regions as it plays a pivotal role in the transmission of functional information due to the high concentration of WM tracts linking the precuneus with other brain areas. Our finding of decreased patient GM and increased functional node strength for the region could be related to the functional compensation effects of AD and merits further investigation ([52, 53]). Additionally, the cingulate gyrus (posterior division), calcarine gyrus, and hippocampus, were important in both sMRI and rs-fMRI. Only the thalamus was relevant to both sMRI and dMRI, and the middle frontal gyrus was the only area important to both rs-fMRI and dMRI. Interestingly, the middle frontal gyrus was a region highlighted in both rs-fMRI and dMRI, but not in the sMRI. This further underlines the importance of integrating different (and complementary) modalities to provide a more complete picture of the complex pathological mechanisms.
5.4 Main contributions and outcomes
We proposed a multimodal and explainable neuroimaging DL model for the classification of amyloid- positive or negative status. The main contribution of this work is a framework that i) successfully integrates volumetric features and connectivity information from sMRI, rs-fMRI, and dMRI data, extracted from one 3D-CNN and two GCNs, respectively, ii) obtains competitive classification performance (ACC=) relative to the state-of-the-art with the advantage of avoiding amyloid PET, iii) provides insight into the brain regions that most contribute to final model predictions using a GBP-based post-hoc analysis. Finally, we evaluated the XAI outcomes by assessing the discriminative power of the selected input features across classes.
With these analyses, we identified multiple brain regions that suffered from the effects of A plaque deposition, resulting in atrophy, functional changes, and WM changes relative to the control class. Furthermore, the analysis identified common regions across modalities, including the precuneus, hippocampus, thalamus, cingulate gyrus, calcarine gyrus, and middle frontal gyrus, strengthing the evidence of their involvement in this pathological process.
Our findings highlight the utility of MRI for studying the effects of A deposition and the importance of integrating complementary information to enable a better understanding of the differences related to amyloid status. However, we acknowledge that a main limitation of this work is the limited number of subjects in the dataset having both A information and all the considered MRI modalities. Further studies are thus needed to evaluate the generalizability of this multimodal framework to larger datasets, also including subjects with opposite A status (e.g., CN pos).
6 Conclusion
In this work, we presented a multimodal and explainable DL-based framework for the classification of amyloid- status, exploiting anatomical and connectivity MRI-based information. The application of GBP enabled the identification of the regions most important to the final model predictions, some of which were common across modalities, (e.g., the precuneus, hippocampus, thalamus, cingulate gyrus, calcarine gyrus, and middle frontal gyrus). Our study demonstrates the potential viability of non-invasive MRI-based detection of amyloid- status involving multimodal data, paving the way for further research in this direction.
Acknowledgments
Data collection and sharing for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904) and DOD ADNI (Department of Defense award number W81XWH-12-2-0012). ADNI is funded by the National Institute on Aging, the National Institute of Biomedical Imaging and Bioengineering, and through generous contributions from the following: AbbVie, Alzheimer’s Association; Alzheimer’s Drug Discovery Foundation; Araclon Biotech; BioClinica, Inc.; Biogen; Bristol-Myers Squibb Company; CereSpir, Inc.; Cogstate; Eisai Inc.; Elan Pharmaceuticals, Inc.; Eli Lilly and Company; EuroImmun; F. Hoffmann-La Roche Ltd and its affiliated company Genentech, Inc.; Fujirebio; GE Healthcare; IXICO Ltd.; Janssen Alzheimer Immunotherapy Research & Development, LLC.; Johnson & Johnson Pharmaceutical Research & Development LLC.; Lumosity; Lundbeck; Merck & Co., Inc.; Meso Scale Diagnostics, LLC.; NeuroRx Research; Neurotrack Technologies; Novartis Pharmaceuticals Corporation; Pfizer Inc.; Piramal Imaging; Servier; Takeda Pharmaceutical Company; and Transition Therapeutics. The Canadian Institutes of Health Research is providing funds to support ADNI clinical sites in Canada. Private sector contributions are facilitated by the Foundation for the National Institutes of Health (www.fnih.org). The grantee organization is the Northern California Institute for Research and Education, and the study is coordinated by the Alzheimer’s Therapeutic Research Institute at the University of Southern California. ADNI data are disseminated by the Laboratory for Neuro Imaging at the University of Southern California.
This study was funded by NIH grant (RF1AG063153) and NSF Grant (2112455), as well as Fondazione CariVerona (EDIPO project, num. 2018.0855.2019) and MIUR D.M. 737/2021 “AI4Health: empowering neurosciences with eXplainable AI methods".
CRediT authors contributions statement
Giorgio Dolci: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Writing - original draft. Charles A. Ellis: Methodology, Writing - original draft. Federica Cruciani: Methodology, Data curation, Writing - review & editing. Lorenza Brusini: Data curation, Writing - review & editing. Anees Abrol: Data curation, Writing - review & editing. Ilaria Boscolo Galazzo: Investigation, Data curation, Writing - review & editing. Gloria Menegaz: Conceptualization, Supervision, Writing - review & editing, Funding acquisition. Vince D. Calhoun: Conceptualization, Supervision, Writing - review & editing, Funding acquisition.
Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Additional information
The data used in this work were collected by ADNI (https://adni.loni.usc.edu/), and they are publically available after requested on the ADNI website.
Ethical standard
The data used in this work were acquired by ADNI (https://adni.loni.usc.edu/). Information regarding the ethical standard and informed consent of ADNI 3 Protocol are available at the following link on the: https://adni.loni.usc.edu/wp-content/themes/freshnews-dev-v2/documents/clinical/ADNI3_Protocol.pdf.

### v2

@@
-Multimodal MRI-based Detection of Amyloid Status in Alzheimer’s Disease Continuum
+Multimodal MRI Accurately Identifies Amyloid Status in Unbalanced Cohorts in Alzheimer’s Disease Continuum
 Abstract
-Amyloid- (A) plaques in conjunction with hyperphosphorylated tau proteins in the form of neurofibrillary tangles are the two neuropathological hallmarks of Alzheimer’s disease (AD).
-In particular, the accumulation of A plaques, as evinced by the A/T/N (amyloid/tau/neurodegeneration) framework, marks the initial stage.
-Thus, the identification of individuals with A positivity could enable early diagnosis and potentially lead to more effective interventions.
-Deep learning methods relying mainly on amyloid PET images have been employed to this end.
-However, PET imaging has some disadvantages, including the need of radiotracers and expensive acquisitions.
-Hence, in this work, we propose a novel multimodal approach that integrates information from structural, functional, and diffusion MRI data to discriminate A status in the AD continuum.
-Our method achieved an accuracy of .
-Furthermore, a post-hoc explainability analysis (guided backpropagation) was performed to retrieve the brain regions that most influenced the model predictions.
-This analysis identified some key regions that were common across modalities, some of which were well-established AD-discriminative biomarkers and related to A deposition, such as the hippocampus, thalamus, precuneus, and cingulate gyrus.
-Hence, our study demonstrates the potential viability of MRI-based characterization of A status, paving the way for further research in this domain.
+Amyloid- (A) plaques in conjunction with hyperphosphorylated tau proteins in the form of neurofibrillary tangles are the two neuropathological hallmarks of Alzheimer’s disease.
+It is well-known that the identification of individuals with A positivity could enable early diagnosis.
+In this work, we aim at capturing the A positivity status in an unbalanced cohort enclosing subjects at different disease stages, exploiting the underlying structural and connectivity disease-induced modulations as revealed by structural, functional, and diffusion MRI.
+Of note, due to the unbalanced cohort, the outcomes may be guided by those factors rather than amyloid accumulation.
+The partial views provided by each modality are integrated in the model allowing to take full advantage of their complementarity in encoding the effects of the A accumulation, leading to an accuracy of .
+The specificity of the information brought by each modality is assessed by post-hoc explainability analysis (guided backpropagation), highlighting the underlying structural and functional changes.
+Noteworthy, well-established biomarker key regions related to A deposition could be identified by all modalities, including the hippocampus, thalamus, precuneus, and cingulate gyrus, witnessing in favor of the reliability of the method as well as its potential in shading light on modality-specific possibly unknown A deposition signatures.
 1 Introduction
@@
 Additionally, A plaques in AD are also linked with demyelination of white matter (WM) tracts ([4]).
-The precise identification of A proteins, and thus the classification of patients as A positive or negative, is only possible through amyloid positron emission tomography (PET) scans and cerebrospinal fluid (CSF) tests.
-Despite their undeniable utility, their use has some negative implications.
-Specifically, amyloid PET uses ionizing radiation and radioactive tracers, is expensive, and is not widely available ([5, 6]).
+In clinical practice, the precise identification of A proteins, and thus the classification of patients as A positive or negative, is performed through amyloid positron emission tomography (PET) scans and cerebrospinal fluid (CSF) tests.
+Despite their undeniable utility, their use has some inherent limitations.
+First, limiting to one modality hides informative features that would require other kinds of investigation tools to be revealed.
+Then, amyloid PET uses ionizing radiation and radioactive tracers, is expensive, and is not widely available ([5, 6]).
 Additionally, the CSF test requires an invasive lumbar puncture.
-In contrast, common, non-invasive MRI techniques can help identifying the effects of A plaques in the brain.
-In particular, structural MRI (sMRI) has been widely used for AD detection and early-prediction due to its ability to detect GM atrophy and structural changes.
+A wider perspective on the disease can be obtained by adding views obtained by other imaging modalities.
+Among these, non-invasive MRI techniques are at the top of the list, allowing to assess structural and functional changes encoding the effect of the amyloid accumulation that can profitably be used as additional biomarkers.
+Structural MRI (sMRI) has been widely used for AD detection and early prediction due to its ability to detect GM atrophy and structural changes.
 Resting-state functional MRI (rs-fMRI) detects changes in blood oxygenation level-dependent (BOLD) signals, which depend on neurovascular coupling, and hence indirectly measures brain neural activity ([7]).
-Different resting-state networks (RSNs), like the saliency and default-mode (DM) networks, are closely related to AD pathology, showing functional alterations ([8, 9, 10]).
-Finally, diffusion MRI (dMRI) is an imaging technique that relies on the movement of water molecules to mark WM boundaries enabling the reconstruction of fiber trajectories and the study of brain structural connections in AD ([4]).
-Additionally, dMRI data are used to extract microstructural information ([11]).
-These two modalities are generally used to map whole-brain connectomes, with rs-fMRI describing the functional connectivity (FC) between regions (usually in terms of correlations) and dMRI the structural connectivity (SC) (most often relying on the number of WM fibers).
-Currently, different studies focusing on A classification tasks (positive vs negative) in AD research have employed PET images and related features with deep learning (DL) models, while only a few works used MRI data.
-[12] considered a slice-level approach for the identification of the A status in conjunction with a 2D-CNN for feature extraction and classification.
-In the same manner, [13] adopted the slice-level approach for the classification of A status employing two different 2D-CNN (ResNet-50 and ResNet-152).
-These two works were able to reach a high level of accuracy (around ) using a slice-level method.
+Different resting-state networks (RSNs), like the saliency and default-mode (DM) networks, have been shown to be altered in AD pathology ([8, 9, 10]).
+Finally, diffusion MRI (dMRI) is an imaging technique that relies on the movement of water molecules, allowing the assessment of both microstructural ([11]) and structural connectivity (SC) changes ([4]).
+These two modalities are generally used to map the whole-brain connectomes, with rs-fMRI describing the functional connectivity (FC) between region pairs (usually in terms of correlations) and dMRI the SC (most often relying on the number of WM fibers linking the target regions).
+Different studies focusing on A classification tasks (positive vs negative) in AD research have employed PET images and related features with deep learning (DL) models, while only a few works used MRI data.
+In [12], a slice-level approach for the identification of the A status was considered in conjunction with a 2D-CNN for feature extraction and classification.
+In the same manner, [13] adopted the slice-level approach for the classification of A status employing two different 2D-CNN (ResNet-50 and ResNet-152), both reaching high accuracy (around ).
 On the other hand, [6] considered three different well-known 3D-CNNs for detecting A positivity relying on 3D Florbetaben brain PET images.
@@
 [15] faced the problem of A classification relying on both early and delay-phase FBB PET images and tested their models considering both single and fused modalities.
-Using FDG and Amyloid PET-derived images, they were able to reach competitive performance in this task, with accuracies around .
+Using FDG and Amyloid PET-derived images, they were able to achieve competitive performance in this task, with accuracies of around .
 More recently, [16] employed FDG-PET-derived features in order to predict the A status in the AD continuum.
 To this end, they tested eight different feature selection methods and eight different classifiers.
-LASSO in conjunction with the Gaussian Naive Bayes (GNB) model performed better with respect to the others, achieving an AUC of 0.924.
+LASSO in conjunction with the Gaussian Naive Bayes (GNB) model performed better with respect to the others, achieving an AUC of .
 Regarding MRI-based approaches for A detection, [17] employed sMRI images in conjunction with a 3D-CNN, while [18] used an SDF-based convolutional network to analyze the hippocampus region.
@@
 Due to the heterogeneous factors that lead to AD, in recent years, many studies have focused on multimodal DL models due to their ability to integrate information of different nature and to outperform single-modality methods ([19, 17, 20]).
-Moreover, among the different DL models, graph neural and convolutional networks (GNNs/GCNs) are becoming popular in neuroscience due to their perfect fit to functional and structural brain networks.
-[21] proposed a benchmark for analyzing fMRI and dMRI networks through GNNs, testing different messages passing, node features, and pooling operations, while Wee and colleagues developed a GCN to study the cortical thickness ([22]).
+Recently, graph neural and convolutional networks (GNNs/GCNs) have become popular in neuroscience due to their perfect fit for functional and structural brain networks.
+[21] proposed a benchmark for analyzing fMRI and dMRI networks through GNNs, testing different messages passing, node features, and pooling operations, while [22] developed a GCN to study the cortical thickness.
 Although DL models can achieve high performance in different tasks, they do not easily provide interpretable output for what they have learned, which is particularly problematic in clinical and biomedical domains.
-To address this issue, eXplainable Artificial Intelligence (XAI) methods have been developed allowing to identify the contributions of input features to final predictions potentially highlighting crucial physiological information for AD ([23, 24, 25, 26]).
-In this study, we present a multimodal and explainable DL framework for the classification of A positive and negative status in the AD continuum.
-Our framework included structural and functional connectomes derived from dMRI and rs-fMRI, respectively, along with sMRI-derived GM 3D volumes in order to investigate complementary aspects as well as their relations.
-This was complemented with an extensive post-hoc XAI analysis to uncover the input features that most influenced the final predictions.
-2 Methods
+To address this issue, eXplainable Artificial Intelligence (XAI) methods have been developed allowing to identify the contributions of input features to final predictions, potentially highlighting crucial information for AD ([23, 24, 25, 26]).
+In this study, the overall goal of the proposed approach consists of decrypting the signatures induced by A accumulation in the views provided by s/rs-f/dMRI, taking advantage of their complementarity for capturing the A status while exploiting their interplay for the identification of the regions that played a prominent role in determining the outcome.
+To instantiate this idea, we present a multimodal and explainable DL framework for the classification of A positive and negative status relying on an unbalanced cohort of individuals spanning the AD continuum.
+The proposed framework includes structural and functional connectomes derived from dMRI and rs-fMRI, respectively, along with sMRI-derived GM 3D volumes in order to investigate complementary aspects as well as their relations.
+For the sake of interpretability of the results, an extensive post-hoc XAI analysis was then performed, pointing the spotlight on the input features that most influenced the final outcome.
+2 Materials and Methods
 2.1 Dataset
@@
 Similarly to [5], a cutoff of was used to define the A status (i.e., for positivity).
-After that, A negative CN, SMC, and EMCI (, , and subjects, respectively) were considered as ’control class’, while A positive EMCI, LMCI, and AD (, , and patients, respectively) formed the ’patient class’, resulting in a final cohort of individuals.
-Table 1 shows the demographic information of the final cohort.
+Aiming at the classification based on the A status as the target outcome, the considered group of individuals was further split gathering the A negative CN, SMC, and EMCI individuals in the NEG class (, , and subjects), and the A positive EMCI, LMCI, and AD in the POS class (, , and subjects), respectively, resulting in an unbalanced data split with respect to the disease stage.
+The LMCI and AD groups were not included in the A negative class since they could represent different underlying conditions linked with the functional decline (e.g., Lewy Body dementia, Frontotemporal dementia, vascular dementia, TDP-43 pathology).
+In contrast, CN and SMC individuals with amyloid accumulation are subjects of ongoing debate, with no clear consensus on whether they represent a prodromal stage of AD or individuals at higher risk of developing the disease but still may never develop the disease ([27, 28]).
 Status # of subjects Age Sex (M/F) MMSE A CN SMC EMCI LMCI AD A - - - A + - -
-The MRI images for the considered cohort were collected as follows: i) T1-weighted sMRI: TR/TE=shortest, TI= ms, FOV= mm2, mm isotropic resolution, slices=; ii) rs-fMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, volumes in almost all subjects, with minimal variations (e.g., ) in a small subset; iii) single-shell dMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, b= and s/mm2.
+The MRI images for the considered cohort were collected as follows: i) T1-weighted sMRI: TE/TR=shortest, TI= ms, FOV= mm2, mm isotropic resolution, slices=; ii) rs-fMRI: TE/TR= ms, FOV= mm3, mm isotropic resolution, volumes in almost all subjects, with minimal variations (e.g., ) in a small subset; iii) single-shell dMRI: TE/TR= ms, FOV= mm3, mm isotropic resolution, b= and s/mm2.
 2.2 Preprocessing and Feature extraction
@@
 In detail, the DL architecture used for the classification of A status (positive/negative) has two modules: i) a feature reduction module that actuates feature reduction using three different Neural Networks (NNs) to transform the input data into corresponding latent representations; and ii) a data fusion & classification module that concatenates the latent representations of each modality and uses a Multilayer Perceptron (MLP) that takes the fused latent features as input for the final classification.
-Finally, a post-hoc explainability analysis was performed on the correctly classified subjects to obtain feature contributions for the classification task.
+Finally, a post-hoc explainability analysis was performed on the correctly classified subjects to highlight the feature contributions to the classification task.
 In the following paragraphs, the two modules are further described.
@@
 Conversely, both FNC and SC graphs were analyzed using two different GCNs.
-These models update the representation of the node aggregating neighbor information, iteratively in each layer, through the message-passing scheme.
+These models update the representation of each node, aggregating neighbor information iteratively in each layer through the message-passing scheme.
 The rs-fMRI channel was analyzed using a GCN with two graph convolutional layers proposed by [32] that were followed by a dense layer.
 This type of convolutional layer is a powerful operator that integrates both high- and low-level structures along with their relationships into a single graph ([32]).
-The dMRI channel was analyzed with a different GCN with two Chebyshev spectral graph convolutional layers proposed by [31] and then followed by a dense layer.
-Spectral convolutional layers are high-performance layers that have been used effectively in irregular graphs ([32]).
+The dMRI channel was analyzed with a different GCN with two Chebyshev spectral graph convolutional layers proposed in [33] followed by a dense layer.
+Spectral convolutional layers are high-performance layers that have been used effectively on irregular graphs ([34]).
 These two architectures were chosen on an empirical basis as those leading to the best unimodal classification accuracy.
@@
 2.4 Training scheme and evaluation
-The model was trained with stratified -fold cross-validation on the entire cohort to investigate its generalizability across individuals, and a hyperparameter search was performed empirically.
-The mini-batch strategy (with subjects per batch) was adopted.
+The model was trained with stratified -fold cross validation on the entire cohort to investigate its generalizability across individuals, and a hyperparameter search was performed empirically through a grid search procedure to maximize the average validation accuracy.
+To this end, different combinations were tested for the hyperparameters (batch size, learning rate, number of epochs, and regularization parameter).
+Additionally, different numbers of hidden layers for CNN, GCNs, and classifier were considered, also changing the number of channels for the convolutional layers and the number of neurons in the dense layers.
+The mini-batch strategy (with subjects per batch) was finally adopted.
 The Adam optimizer (learning rate: ) was used to update the entire multimodal architecture.
@@
 Performance was evaluated using the mean evaluation accuracy, precision, recall, and F1 score over the five folds.
-For comparison, we also tested the unimodal models with the corresponding MRI data, where the architectures were the same of the different branches of the multimodal framework (i.e., 3D-CNN for sMRI, and GCN for both rs-fMRI and dMRI).
+For the sake of comparison, we also tested the unimodal models with the corresponding MRI data, where the architectures were the same as the different branches of the multimodal framework (i.e., 3D-CNN for sMRI, and GCN for both rs-fMRI and dMRI).
 3 Post-hoc analysis
@@
 3.2 Contribution maps and Statistical analysis
-The attribution maps were extracted for the correctly classified A positive subjects, and in particular, we focused on the average A positive attribution map to identify the most important features.
+The attribution maps were extracted for the correctly classified A positive subjects.
+The average A positive subject attribution map was derived for identifying the most important features.
 To evaluate the sMRI GBP contribution maps, the Harvard-Oxford ([31]) and the probabilistic cerebellar ([37]) atlases from FSL were employed to define different ROIs, including cortical, subcortical, and cerebellum regions.
@@
 Conversely, the GBP attribution of each node for both rs-fMRI and dMRI was extracted directly from the two GCNs.
-Then, considering the average map for each modality, the percentage of explanation for each region/node was computed over the total contribution within and between each modality.
+Then, considering the average map for each modality, the percentage of explanation for each region/node was computed over the total contribution within and across modalities.
 The top ROIs (sMRI) and nodes (rs-fMRI and dMRI) with the highest percentage of GBP contribution were selected for further investigation.
@@
 For the sMRI, the mean values of the top ROIs resulting from the XAI analysis were extracted from the input GM volumes and used as features for the statistical analysis.
-For the rs-fMRI and dMRI, graph-based measures were first derived from the full connectivity matrices in order to have a summary measure per node, and then we considered only the top nodes for both rs-fMRI and dMRI for the statistical analysis.
-In particular, the node strength was computed for rs-fMRI and FNC matrices.
+For the rs-fMRI and dMRI, graph-based measures were first derived from the full connectivity matrices in order to have a summary measure per node, and then only the top nodes were retained for both rs-fMRI and dMRI for statistical analysis.
+In particular, the node strength was computed for the FNC matrices.
 This is defined as the sum of the weights of the edges connected with a given node, where in absolute terms, higher values mean more important nodes.
@@
 These two metrics were computed using the Brain Connectivity Toolbox (BCT) ([38]) in Matlab.
-Mann-Whitney tests were then used to compare the values of all these features for the top ROIs/nodes between A positive and negative correctly classified patients.
-Finally, FDR correction for multiple comparisons was applied to the p-values obtained from the statistical tests.
+Mann-Whitney tests were then used to compare the values of all these features for the top ROIs/nodes between A positive and negative correctly classified individuals.
+Finally, FDR correction for multiple comparisons was applied.
 4 Results
@@
 The single networks that composed the multimodal framework along with the corresponding input data were also tested in the same classification task.
-Table 2 shows the performance comparisons for the multimodal method and the three unimodal models.
-Results highlight how the multimodal framework is able to outperform the unimodal models in terms of accuracy, recall, and F1 score in the same classification task.
+Table 2 shows the performance comparisons for the multimodal and the three unimodal models.
+Results highlight how the multimodal pipeline is able to outperform the unimodal models in terms of accuracy, recall, and F1 score in the same classification task.
 Model ACC PRE REC F1 Multimodal Unimodal sMRI Unimodal fMRI Unimodal dMRI
 4.2 GBP-based attribution maps
-The evaluation set in the fold with the highest evaluation accuracy was used for the post-hoc analysis. Importantly, in the analyzed fold, the evaluation subjects were not used for training.
-Figure 2A shows the sMRI GBP attribution map for the mean A positive patient overlaid to the MNI template.
-Qualitatively, mainly subcortical regions (e.g., hippocampus, thalamus) are relevant to the final classification along with a few cortical areas.
+The evaluation set in the fold with the highest evaluation accuracy was used for post-hoc analysis. Importantly, in the analyzed fold, the evaluation and training sets were disjoint.
+Figure 2A shows the sMRI GBP attribution map for the A mean positive subject, overlaid to the MNI template.
+Qualitatively, mainly subcortical regions (e.g., hippocampus, thalamus) were relevant to the final classification along with a few cortical areas.
 Figure 2B shows the most important nodes for the rs-fMRI modality.
@@
 The most important nodes belong to the DM ( nodes), CC ( nodes), VI ( nodes), and SM ( node) networks.
-Finally, Figure 2C shows the most important dMRI nodes (ROIs). Subcortical regions (thalamus, cerebellum) along with areas in the frontal, temporal, and parietal lobes are identified.
+Finally, Figure 2C shows the most important dMRI nodes (ROIs). Subcortical regions (thalamus, cerebellum) along with areas in the frontal, temporal, and parietal lobes were identified.
 4.3 Feature relevance & Statistical analysis
-Figure 3A shows violinplots for the top sMRI features, while Table 3a shows the corresponding percentages of mean GBP contribution across A positive correctly classified patients for the same regions, weighted by their volumes.
+Figure 3A shows the violinplots for the top sMRI features, while Table 3a shows the corresponding percentages of mean GBP contributions across A positive correctly classified subjects for the same regions, weighted by their volumes.
 The Table also reports the p-values and FDR-corrected p-values of the statistical tests performed on the input features.
 We identified significant differences between classes after FDR correction in the precuneus, cingulate gyrus (posterior division), thalamus, hippocampus, supracalcarine and intracalcarine cortices, amygdala, and temporal occipital fusiform cortex showing an overall comparison direction of .
-Figure 3B shows violinplots representing the graph node strength for the top rs-fMRI nodes, while Table 3b shows the mean percentages of GBP contribution extracted from the same nodes (ICs) across A positive patients.
+Figure 3B shows violinplots representing the graph node strength for the top rs-fMRI nodes, while Table 3b shows the mean percentages of GBP contributions extracted from the same nodes (ICs) across A positive patients.
 After FDR correction, the precuneus (IC48) was the only region with significant differences ().
-Before correction, another precuneus component (IC43) and precentral gyrus (IC14) were also significant with , while the left inferior parietal lobule (IC38) exhibited a trend towards uncorrected statistical significance (p, with ).
-Figure 3C shows violinplots representing the graph betweenness centrality for the top dMRI nodes, and Table 3c shows the mean percentages of GBP contribution for the same nodes across A positive correctly classified patients.
-The precuneus LH and cerebellum RH were significant when considering uncorrected p-values (), but none of the results survived multiple comparison corrections.
+Before correction, precuneus component (IC43) and precentral gyrus (IC14) were also significant with , while the left inferior parietal lobule (IC38) exhibited a trend towards uncorrected statistical significance (p, with ).
+Figure 3C shows the violinplots representing the graph betweenness centrality for the top dMRI nodes, and Table 3c shows the mean percentages of GBP contributions for the same nodes across A positive correctly classified subject.
+The precuneus LH and cerebellum RH were significant when considering uncorrected p-values (), but none of the results survived multiple comparisons correction.
 Moreover, the right superior parietal gyrus showed a trend towards significance (p, ).
 5 Discussion
-In this work, we presented a multimodal data fusion framework that integrates multiple MRI techniques (sMRI, rs-fMRI, and dMRI) for the classification of A status.
-A post-hoc XAI analysis and evaluation was performed using GBP to identify feature importance, complemented by statistical analyses on the identified input features for validation via post-hoc plausibility assessment.
-The underlying assumption is that maximally contributing regions would bring discriminative information across the target classes.
+In this work, we proposed a multimodal data fusion framework that integrates multiple MRI techniques (sMRI, rs-fMRI, and dMRI) for the classification of A status in an unbalanced cohort of subjects.
+A post-hoc XAI analysis and evaluation was performed using GBP to assess feature importance, complemented by statistical analyses on the so-identified input features for validation via post-hoc plausibility assessment.
 a) Structural MRI
@@
 5.1 Classification performance
-Our results showed that the multimodal framework outperformed single-modality models in terms of classification performance, in particular for rs-fMRI and dMRI, while the sMRI achieved performance close to the multimodal one, but with high variance across the evaluation folds.
-This confirms the added value provided by multimodality approaches which improve classification accuracy and, at the same time, carry more information, particularly relevant for the study of complex neurodegenerative diseases influenced by multi-domain factors.
-Table 4 shows the performance of state-of-the-art works for the classification of A positive and negative subjects.
-Of note, current state-of-the-art works use multiple datasets and approaches for addressing this task without considering a common benchmark dataset.
-Hence, the comparison of the different approaches is imperfect and any resulting conclusions held tentatively.
-Most of the works relied on PET scans to address this task, e.g., [14] developed a CNN model that takes in input the different views (axial, coronal, and sagittal) of a 3D volume; while [6] used three different well-known 3D-CNN (Inception3D, ResNet3D, and VGG3D) to address this task using the full 3D volume of PET scans as input.
+Results showed that the multimodal framework outperformed single-modality models in terms of classification performance, in particular for rs-fMRI and dMRI, while the sMRI achieved performance close to the multimodal one, but with higher variance across the evaluation folds.
+This provides evidence of the added value brought by multimodality approaches in terms of classification accuracy while injecting complementary information shading light on the underlying neurophysiological mechanisms, which is particularly relevant for the study of complex neurodegenerative diseases influenced by multi-domain factors.
+Table 4 shows the performance of state-of-the-art (SOA) works for the classification of A positive versus negative subjects.
+Of note, current SOA works use multiple datasets and approaches for addressing this task.
+It goes without saying that the lack of a common reference dataset inherently limits the relevance of performance comparison.
+However, reaching good accuracy and outperforming the SOA does not exhaust the contribution of the proposed approach, whose potential lies in decoding the neurophysiological changes induced by the A status as captured by the considered MRI modalities.
+Most of the SOA works relied on PET scans to address this task, e.g., [14] developed a CNN model that takes the different views (axial, coronal, and sagittal) of a 3D volume as input, while [6] used three different well-known 3D-CNN (Inception3D, ResNet3D, and VGG3D) models to address this task using the full 3D volume of PET scans.
 They achieved an accuracy of around on average in the test set, with a maximum accuracy of by [6] using the VGG3D architecture.
@@
 They achieved an AUC of employing LASSO with the GNB model.
-On the other hand, only few works have used MRI modalities to classify A positive and negative subjects.
-[18] used only the hippocampus region for predicting A positivity, testing their network on two different classification tasks: AD A POS versus CN A NEG, and MCI A POS versus MCI A NEG.
-They achieved an accuracy of and , in the first and second tasks, respectively.
-Lastly, [17] implemented a 3D-CNN for detecting A status from sMRI images defining the A status of the CN, MCI, and AD patients considering the PET cortical standardized uptake value ratio.
-They achieved an accuracy of .
-Comparing our proposed method with these five works, we outperformed the method that relied on 3D FDG-PET proposed by [14] and the performance of [17].
-Our study achieved competitive performance in the validation set compared to [6] using amyloid PET images, although our framework achieved a higher accuracy compared to the ResNet3D they used, and a similar accuracy to the one reached by the model proposed by [18] using the hippocampus region only.
-While the current classification of A positive or negative subjects can only be performed through CSF tests or amyloid PET scans with high precision, the development of different approaches based on non-invasive MRI modalities for determining A status could improve patient care highlighting specific mechanisms that underlie AD.
-We further showed the ability of our model to successfully integrate structural GM features extracted from a 3D-CNN along with functional and structural connectivity analyzed by two types of GCNs.
-Notably, while a few studies have recently started to explore SC/FC in conjunction with GCNs in the AD classification tasks ([38, 39]), results are still limited in this specific problem calling for further investigation in the AD continuum.
+Only a few works used MRI modalities to classify A positive and negative subjects.
+[18] focused on the hippocampus region only for detecting A positivity, testing the network on two different classification tasks: AD A positive versus CN A negative, and MCI A positive versus MCI A negative, achieving an accuracy of and , in the first and second tasks, respectively.
+Lastly, [17] implemented a 3D-CNN for detecting A status from sMRI images defining the A status of the CN, MCI, and AD patients considering the PET cortical standardized uptake value ratio reaching an accuracy of .
+Comparing the proposed method with these six works, we outperformed the method that relied on 3D FDG-PET proposed by [14] and the performance of [17].
+Our study achieved competitive performance in the validation set compared to [6] using amyloid PET images, although our pipeline led to higher accuracy compared to the ResNet3D they used, and a similar accuracy to that in [18] using the hippocampus region only.
+Besides providing competitive performance, our model allows to successfully integrate structural GM features extracted from a 3D-CNN along with functional and structural connectivity relying on two ad-hoc GCNs.
+Notably, while a few studies have recently started to explore SC/FC in conjunction with GCNs in the AD classification task ([40, 41]), results are still limited, calling for further investigation in the AD continuum.
 Authors Modalities Study cohort Input data Model ACC REC PRE F1 [14] A/FDG-PET POS, NEG 3D FDG-PET 2.5-D CNN - - n.d.
@@
 5.2 Explainability analysis
-The post-hoc analysis, performed on the correctly classified subjects, consisted of two steps: i) extraction of GBP attribution maps for A positive patients and corresponding percentage of contribution for each brain region both within and between modalities for the average positive patient; ii) statistical analysis on the input features of the most important regions/nodes identified by GBP.
-In our analysis of sMRI attribution maps, multiple brain regions involved in AD neurodegeneration were identified.
-These regions had a higher percentage of contribution relative to the others.
-These regions belonged to subcortical areas (hippocampus, thalamus, putamen, accumbens, and amygdala) and temporal/occipital/parietal areas (posterior cingulate gyrus, precuneus cortex, supracalcarine and intracalcarine cortices, and the temporal occipital fusiform cortex).
+The post-hoc analysis, performed on the correctly classified subjects, consisted of two steps: i) extraction of GBP attribution maps for A mean positive subjects as well as the corresponding percentage of contribution for each brain region, both within and between modalities; ii) statistical analysis on the input features of the most important regions/nodes identified by GBP.
+The analysis of the sMRI attribution maps revealed that multiple brain regions involved in AD neurodegeneration could be identified.
+These regions had a higher percentage of contribution relative to the others and belonged to subcortical areas (hippocampus, thalamus, putamen, accumbens, and amygdala) and temporal/occipital/parietal areas (posterior cingulate gyrus, precuneus cortex, supracalcarine and intracalcarine cortices, and the temporal occipital fusiform cortex).
 The hippocampus, in particular, is a well-known biomarker for AD that is subject to high levels of atrophy.
@@
 An important region that is connected with the hippocampus is the cingulate gyrus which showed a strong reduction of GM in AD patients ([43, 44]).
-In a clinical study performed by [43], significant associations were detected between A accumulation and GM atrophy in the hippocampus and posterior cingulate gyrus for the MCI and AD A+ patients.
-Moreover, other subcortical structures, like the thalamus and putamen, have a high atrophy rate in clinical patients relative to healthy individuals ([44, 43]), probably due to A deposition.
-A previous study highlighted an increase in SUVR (derived from florbetapir PET amyloid) in both putamen and thalamus in the preclinical stages of AD ([45]).
-In the statistical analysis performed on the input GM volumes, significant differences were detected in the hippocampus, posterior cingulate gyrus, amygdala, thalamus, and precuneus, which demonstrated a considerable decrease of GM in the A positive patients.
+In a clinical study performed by [45], significant associations were detected between A accumulation and GM atrophy in the hippocampus and posterior cingulate gyrus for the MCI and AD A positive subjects.
+Some works also linked the nucleus accumbens region to AD pathology and progression.
+Evidence of GM loss and alteration of nucleus accumbens in MCI and AD, with respect to the CN, had previously been highlighted ([46, 47]).
+Additionally, [48] showed how A oligomers in nucleus accumbens can promote synaptic loss and motivation deficits in AD.
+Other subcortical structures, like the thalamus and putamen, have a high atrophy rate in clinical patients relative to healthy individuals ([49, 45]), probably due to A deposition.
+A previous study highlighted an increase in the standardized uptake value ratio (derived from florbetapir PET) in both putamen and thalamus in the preclinical stages of AD ([50]).
+Following the statistical analysis performed on the input GM volumes, significant differences were detected in the hippocampus, posterior cingulate gyrus, amygdala, thalamus, and precuneus, highlighting a considerable decrease of GM in the A positive subjects.
 These findings were consistent with preexisting literature.
@@
 For the rs-fMRI, the brain regions (ICs) that most contributed to the final classification resided in the DM and CC networks, along with three regions for VI and SM networks (two and one, respectively).
-The DM network is involved in memory, self-knowledge, and thinking, and it is highly related to AD ([8, 41]) Previous studies also suggested that DMN is associated with A plaques ([47]).
+The DM network is involved in memory, self-knowledge, and thinking, and it is highly related to AD ([8, 43]).
+Previous studies also suggested that DMN is associated with the presence of A plaques ([52]).
 The CC network is associated with selective attention, working memory, and stimulus-response mapping ([53, 54]).
-The statistical analysis, performed on the node strength values, detected one region with significant differences between the two groups after FDR correction and three regions before (uncorrected p-values).
+The statistical analysis, performed on the node strength values, detected one region with a significant difference between the two groups after FDR correction and three regions before (uncorrected p-values).
 One more region, the left inferior parietal lobule, exhibited a trend toward significance.
@@
 Previous works also identified the posterior DM (the precuneus and posterior cingulate cortex) ([9]) and parts of the CC network (i.e., the middle frontal gyrus and hippocampus) ([59, 60]) as strongly affected by A deposition.
-Lastly, in the dMRI channel, GBP identified multiple cortical and subcortical regions as relevant to the final predictions of our model.
-The cerebellum was also revealed to be important by the model.
+Lastly, in the dMRI channel, GBP identified multiple cortical and subcortical regions as relevant to the outcome of our model.
+The cerebellum resulted to be important for the model.
 Of note, the rostral middle frontal gyrus was assigned a high percentage of contribution relative to the other regions.
@@
 After FDR correction, no region had statistically significant group differences in betweenness centrality values.
-On the other hand, based upon uncorrected p-values, significant differences (p) were highlighted in the precuneus and cerebellum (right hemisphere), while the superior parietal region was close to significance (p=).
+On the other hand, based upon uncorrected p-values, significant differences (p) were detected in the precuneus and cerebellum (right hemisphere), while the superior parietal region was close to significance (p=).
 Regarding the two nodes with significant uncorrected differences, the precuneus, which is also identified in the rs-fMRI, is an important hub for functional operations because it is highly connected with other regions by both short- and long-range WM fibers ([58]).
@@
 Widespread structural alterations in WM tracts connecting cortical and subcortical regions in both hemispheres were detected between A negative CN and positive preclinical AD patients by [64].
-Finally, when combining the explanations of all channels, we found that sMRI most contributed to the final prediction.
+When combining the explanations of all channels, we found that sMRI most contributed to the final prediction.
 This was expected as we initially assumed that the effects of A plaques would predominantly impact GM volumes.
@@
 It was one of the most interesting regions as it plays a pivotal role in the transmission of functional information due to the high concentration of WM tracts linking the precuneus with other brain areas.
-Our finding of decreased patient GM and increased functional node strength for the region could be related to the functional compensation effects of AD and merits further investigation ([52, 53]).
-Additionally, the cingulate gyrus (posterior division), calcarine gyrus, and hippocampus, were important in both sMRI and rs-fMRI.
+Our finding of decreased GM and increased functional node strength for the region in A positive subjects could be related to the functional compensation effects of AD and deserves further investigation ([57, 58]).
+Additionally, the cingulate gyrus (posterior division), calcarine gyrus, and hippocampus were important in both sMRI and rs-fMRI.
 Only the thalamus was relevant to both sMRI and dMRI, and the middle frontal gyrus was the only area important to both rs-fMRI and dMRI.
 Interestingly, the middle frontal gyrus was a region highlighted in both rs-fMRI and dMRI, but not in the sMRI.
-This further underlines the importance of integrating different (and complementary) modalities to provide a more complete picture of the complex pathological mechanisms.
+This further underlines the importance of integrating different (and complementary) modalities to provide a complete picture of the complex pathological mechanisms.
 5.4 Main contributions and outcomes
 We proposed a multimodal and explainable neuroimaging DL model for the classification of amyloid- positive or negative status.
-The main contribution of this work is a framework that i) successfully integrates volumetric features and connectivity information from sMRI, rs-fMRI, and dMRI data, extracted from one 3D-CNN and two GCNs, respectively, ii) obtains competitive classification performance (ACC=) relative to the state-of-the-art with the advantage of avoiding amyloid PET, iii) provides insight into the brain regions that most contribute to final model predictions using a GBP-based post-hoc analysis.
+The main contribution of this work is a framework that i) successfully integrates volumetric features and connectivity information from sMRI, rs-fMRI, and dMRI data, extracted from one 3D-CNN and two GCNs, respectively, ii) obtains competitive classification performance (ACC=) relative to the state-of-the-art, iii) provides insight into the brain regions that most contribute to final model outcome using a GBP-based post-hoc analysis.
 Finally, we evaluated the XAI outcomes by assessing the discriminative power of the selected input features across classes.
-With these analyses, we identified multiple brain regions that suffered from the effects of A plaque deposition, resulting in atrophy, functional changes, and WM changes relative to the control class.
+With these analyses, we identified multiple brain regions that could be altered by A plaque deposition, resulting in atrophy, functional changes, and WM connectivity changes relative to the negative class.
 Furthermore, the analysis identified common regions across modalities, including the precuneus, hippocampus, thalamus, cingulate gyrus, calcarine gyrus, and middle frontal gyrus, strengthing the evidence of their involvement in this pathological process.
-Our findings highlight the utility of MRI for studying the effects of A deposition and the importance of integrating complementary information to enable a better understanding of the differences related to amyloid status.
-However, we acknowledge that a main limitation of this work is the limited number of subjects in the dataset having both A information and all the considered MRI modalities.
-Further studies are thus needed to evaluate the generalizability of this multimodal framework to larger datasets, also including subjects with opposite A status (e.g., CN pos).
+Our findings highlight the utility of MRI for studying the possible effects of A deposition and the importance of integrating complementary information to enable a better understanding of the differences related to amyloid status.
+However, we acknowledge that the two main limitations of this work are the limited number of subjects in the dataset having both A information and all the considered MRI modalities, and the unbalanced cohort including CN, SMC, and EMCI in the A negative group and EMCI, LMCI, and AD in the A positive one.
+Hence, the outcomes could be guided by those factors that were influenced by the amyloid accumulation as captured by the specific modality.
+In consequence, the GM branch captured atrophy, while the structural and functional connectivity branches captured the respective amyloid-induced connectivity modulations, leading to a combined effect due to A and dementia-related processes.
+Further studies are thus needed to assess the generalizability of the proposed multimodal framework to larger and possibly employing balanced datasets, relying on study cohorts composed of all clinical classes (CN, SMC, EMCI, LMCI, and AD) in both groups analyzed, A negative and positive individuals, in order to disentangle the contributions of the different factors and hence having the outcomes only related to A.
 6 Conclusion